In a group of five friends, the sums of the ages of each group of four of them are 124, 128, 130, 136, and 142. What is the age of the youngest?
Answer by Edwin McCravy(20086)

In a group of five friends, the sums of the ages of each group of four of them are 124, 128, 130, 136, and 142. What is the age of the youngest?

Let the five ages be A, B, C, D, and E.

A + B + C + D = 124
A + B + C + E = 128
A + B + D + E = 130
A + C + D + E = 136
B + C + D + E = 142

We consider that as this system, putting in all the 
understood 1's and 0's

1A + 1B + 1C + 1D + 0E = 124
1A + 1B + 1C + 0D + 1E = 128
1A + 1B + 0C + 1D + 1E = 130
1A + 0B + 1C + 1D + 1E = 136
0A + 1B + 1C + 1D + 1E = 142

Now we make the augmented matrix:



Using my TI-84 I get for the rref of that matrix:



A=23, B=29, C=35, D=37, E=41.

The youngest age is A, which is 23. 

Edwin